OpenStudy (anonymous):

X^2-y=3 x-y=-3 solution set HELP PLEASE! I need a answer not a explanation!

OpenStudy (anonymous):

Look at the question again, I typed it incorrectly the first time

hero (hero):

Now, expand and solve the quadratic equation.

OpenStudy (anonymous):

I need a little more explanation please. This time with the equation X^2-y=3 x-y=-3

hero (hero):

I think you switched things up on me. isolate y in both equations.

hero (hero):

x^2 - 3 = y x + 3 = y Then set y = y x^2 - 3 = x + 3 Then setup a quadratic equation.

OpenStudy (anonymous):

so it would be x^2-3=o and x+3=0?

hero (hero):

No, you skipped a step

hero (hero):

You have to put the equation x^2 - 3 = x + 3 in the form ax^2 + bx + c = 0

OpenStudy (anonymous):

how do I know what goes where?

OpenStudy (anonymous):

I am just now learning this stuff